[Ncep.nhc.nco_contacts] TSB Morning Rounds - Wed Jul 19, 2017
pablo.santos at noaa.gov
Thu Jul 20 12:26:20 UTC 2017
Tickets were submitted on the satellite data ingest issues and the ATCF
this morning for immediate action.
On 7/19/17 8:57 PM, Mark DeMaria - NOAA Federal wrote:
> Thanks for looking in to this. I think there are multiple problems as
> you summarized in your message.
> First, our local GOES-W is not working and we should put in an NCO
> helpdesk ticket for that in the morning as you mentioned. It might be
> fixed just by starting and stopped the global imaging software.
> Second, the problem with the connection to NESDIS might be due to a
> problem there. Early this evening I checked the SAB TC-centered
> imagery on their web page and that data was missing starting right at
> 18 UTC. I am nearly certain that imagery comes from the same NESDIS
> servers that we get our data from.
> The various file system problems are harder to understand but might be
> a result of hung processes on our satellite server. Pablo can
> coordinate a reboot of sat01 with HSU, TAFB and NCO.
> On Wed, Jul 19, 2017 at 5:39 PM, Monica Bozeman - NOAA Federal
> <monica.bozeman at noaa.gov <mailto:monica.bozeman at noaa.gov>> wrote:
> Hi All,
> Forecasters reported that the GOES-W data was out for only the
> MERCATOR and SAT_PROJ directories from 18-20Z today. Upon
> investigation, it seems that our GOES-W data from the local dish
> is out AND we lost the connection to NESDIS during this period for
> reasons unknown. There are also some bizarre things going on on
> that server with writing to the filesystem, so TSB may initiate a
> reboot of the processing server tomorrow before lunchtime to
> rectify the situation. TSB will also be submitting a helpdesk
> ticket for the GOES-W dish tomorrow as well.
> Please let TSB know if there are additional problems with the
> GOES-W data overnight.
> Monica Bozeman
> National Hurricane Center
> 301-683-1553 <tel:%28301%29%20683-1553>
> monica.bozeman at noaa.gov <mailto:monica.bozeman at noaa.gov>
> On Wed, Jul 19, 2017 at 1:34 PM, TSB Admin <nhc.tsbadmin at noaa.gov
> <mailto:nhc.tsbadmin at noaa.gov>> wrote:
> Here is today's summary of NHC computer operations:
> --- Pablo/Mello
> The MIM deadline is now 0830Z to cover folks who release the
> MIM prior to 6Z.
> SDM called after submitting guidance for Fernanda both in
> morning after 12Z submission and after 18Z submission, as they
> didn't receive it. Resubmitted it and they received it.
> Noticed both times that ATCF showed notice ("Warning: Two
> storms with priority = 2"), which appears to be Eugene, even
> though that system is not active. Mike "ended" all the old
> storms in the ATCF that are no longer active. That should take
> care of the
> warning about two storms with the same priority.
> Dan reported this morning that he also got the red banners.
> Spoke to Mike about it and decided to have Monica look into it.
> *Other items:*
> Just a reminder that the operational GFS/GDAS model is
> scheduled to be upgraded today with the 12Z cycle. Mello will
> monitor this today.
> *NHC RFCs:*
> New RFC submitted today for NetApp migration next week and
> discussed in NCO daily call.
Dr. Pablo Santos
Technology and Science Branch
National Hurricane Center, Miami, FL
Twitter: @NWSNHC @NHC_Atlantic @NHC_Pacific @NHC_TAFB @NHC_Surge
-------------- next part --------------
An HTML attachment was scrubbed...
More information about the Ncep.nhc.nco_contacts